OpenStudy (anonymous):

I have a horrid headache over this problem! please help :( ----> 1/6y^2=-y -3/2

OpenStudy (anonymous):

1/6y^2=-y-3/2 y^2=-6y-9 y^2+6y+9=0 (y+3)^2=0 y=-3

OpenStudy (anonymous):

*huggles* Thank you!!!! :D